OpenStudy (anonymous):

Sole by the quadratic formula: 4x^2-8x+1=0 Please explain

OpenStudy (anonymous):

- b + - sqrt(b^2 - 4ac)/(2a) <-----quadratic formula

OpenStudy (anonymous):

b = -8, a = 4, c =1 ---- plug in values and solve

OpenStudy (anonymous):

x=\[[-(-8)+/-\sqrt{(-8^2)-4(4)(1)}]/2(4)\] x=\[{8+/-\sqrt{64-16}}/8\] x=\[8+/-\sqrt{48}/8\] x=\[[8+/-(4\sqrt{3})]/8\] x=[2+\[\sqrt{3}/2\]], 2-\[\sqrt{3}/2\] x=1.866,0.134